OpenStudy (anonymous):

for the function g(x)=x^2-3x+1, find 3g(1)

OpenStudy (anonymous):

You must not be very bright. Just kidding. g(x)=blah g(1)=1*1-3*1+1 = -1 3*(-1)=-3

OpenStudy (anonymous):

oh you multiply by 3 after heh thanks